OpenStudy (anonymous):

(please just get straight to the point, i have a deadline for this assignment and need answers) A marble or small sphere, made of stone, is found to have a mass of 3.05 grams. To find its volume, it is placed in a graduated cylinder that originally contains 10.00 mL of water. When the marble is placed in the graduated cylinder, the new volume of the water and marble together is 11.20 mL. a. Calculate the volume of the marble (Be sure to include units). b. Calculate the density of the marble (Be sure to include units).

OpenStudy (taramgrant0543664):

Sorry I don't have a calculator with me! a) would be the difference between the initial and final volume so Vf-Vi= Vtot which is your answer. Units would be in mL b) to find density you use the formula D=M/V where D is density, V is the volume you calculated in a and M is the mass of the marble that is given. Units for this would be g/mL

OpenStudy (sweetburger):

the marble displaced 1.2mL so that would be the volume of the marble so to find the density being D=M/V so 3.05g/1.2mL = Xg/mL I will let you solve for x which is the density.